OpenStudy (anonymous):

What is the correct classification of the system of equations below? y=3x-2 & y-3x=3

OpenStudy (texaschic101):

(In y = mx + b form, the number in the m position is the slope and the number in the b position is the y intercept.) y = 3x - 2 -- slope = 3 and y intercept = -2 y - 3x = 3 -- put in y = mx + b form y = 3x - 3 -- slope = 3 and y intercept = -3 when you have a system of equations where the slopes are the same but the y intercepts are different, then this system is parallel and has no solutions because the lines never intersect.